HelpMePlz:

What is 4.86 × 104 written in standard form? A. 0.0000486 B. 0.000486 C. 48,600 D. 486,000

Champion:

4.86 × 104 = 505.44 Your question is probably 4.86 × 10^4=\(4.86 × 10^4\) If the exponent of 10 is positive, move the decimal that number of places to the \(right\). Otherwise move to the left. Example: \(2.5\times 10^2=2.5\times 100=250\)